OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (1, -3) and has a slope of 2? (4 points) Answer 2x - y = 5 2x - y = -5 2x + y = 5 2x + y = -5

OpenStudy (anonymous):

2x-y=5

OpenStudy (anonymous):

2x + y = 5

OpenStudy (anonymous):

2x-y=5

OpenStudy (anonymous):

I'm off my game tonight.

OpenStudy (anonymous):

no worries